OpenStudy (anonymous):

Simplify spuare root 84

OpenStudy (anonymous):

will be a surd

OpenStudy (dape):

First factor 84 as much as possible \[ 84 = 4*21 = 2*2*3*7 \] If we look at the square root we have \[ \sqrt{84} = \sqrt{2*2*3*7} = \sqrt{2^2*3*7} = 2\sqrt{3*7} = 2\sqrt{21} \]